OpenStudy (lizilizi28):

quadratic formula y=x^2+6x-5

OpenStudy (anonymous):

wht u wanna find

OpenStudy (lizilizi28):

use the quadratic formula to solve to find any x intercept of the graph equation y=x^2+6x-5

OpenStudy (lizilizi28):

u know

hartnn (hartnn):

to find x intercept, put y=0 so your quadratic equation will be x^2+6x-5 =0 Compare this quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)